Effect of Doubtful Accounts on Net Income During its first year of operations, Master Plumbing Supply Co. had sales of $410,000, wrote off $6,600 of accounts as uncollectible using the direct write-off method, and reported net income of $45,100. Determine what the net income would have been if the allowance method had been used, and the company estimated that 1 1/2% of sales would be uncollectible.
Answer #1
Ans. Net income = Net income (using direct write off method) + Already wrote off amount - Uncollectible under allowance method
$45,100 + $6,600 - $6,150
$45,550
*Calculations for uncollectible amount under allowance method:
Uncollectible rate   =   1   1/2 %   =   (2*1+1) / 2%    =   3 / 2% = 3 / 2 / 100   =   0.015
Uncollectible amount   =   Sales * Uncollectible rate
$410,000 * 0.015
$6,150